b) The life of an electronic device is known to have the exponential distribution

with parameter  = 1

1000 .

(i) What is the probability that the device lasts more than 1000 hours?

[2 marks]

(ii) What is the probability it will last less than 1200 hours?

[2 marks]

(iii) Find the mean and variance of the life of the electronic device.

[4 marks]


1
Expert's answer
2021-09-10T09:58:59-0400

Let "X=" the life of an electronic device: "X\sim Po(\lambda)."

Given "\lambda=1000."

For sufficiently large values of "\u03bb," (say "\u03bb>1,000" ), the "Normal(\u03bc = \u03bb,\u03c3^2 = \u03bb)" Distribution is an excellent approximation to the Poisson"(\u03bb)" Distribution.

(i)


"P(X>1000)=1-P(X\leq 100)"

"\approx1-P(Z\leq\dfrac{1000+0.5-1000}{\sqrt{1000}})"

"\approx1-P(Z\leq 0.01581139)\approx0.49369243"

(ii)


"P(X<1200)=P(Z<\dfrac{1200-0.5-1000}{\sqrt{1000}})"

"\approx P(Z<6.30874393)"

"\approx0.99999999986\approx1"

(iii)


"\mu=\lambda=1000"

"Var(X)=\sigma^2=\lambda=1000"